### v1

# Diffusion-Based Scenario Tree Generation for Multivariate Time Series Prediction and Multistage Stochastic Optimization

Abstract
Stochastic forecasting is critical for efficient decision-making in uncertain systems, such as energy markets and finance, where estimating the full distribution of future scenarios is essential. We propose Diffusion Scenario Tree (DST), a general framework for constructing scenario trees for multivariate prediction tasks using diffusion-based probabilistic forecasting models. DST recursively samples future trajectories and organizes them into a tree via clustering, ensuring non-anticipativity (decisions depending only on observed history) at each stage. We evaluate the framework on the optimization task of energy arbitrage in New York State’s day-ahead electricity market. Experimental results show that our approach consistently outperforms the same optimization algorithms that use scenario trees from more conventional models and Model-Free Reinforcement Learning baselines. Furthermore, using DST for stochastic optimization yields more efficient decision policies, achieving higher performance by better handling uncertainty than deterministic and stochastic MPC variants using the same diffusion-based forecaster [1].
Index Terms—  Diffusion Models, Scenario Trees, Stochastic Optimization, Energy Markets, Uncertainty Quantification
1 Introduction
Robust decision-making under uncertainty requires forecasts that capture not only the most likely outcome but also the full distribution of future evolution. Scenario trees, a classical tool in stochastic optimization, model time-evolving uncertainty while enforcing non-anticipativity, ensuring decisions at each node depend only on information revealed up to that point [2, 3]. Traditional methods like VAR and LSTM are widely used but struggle with complex, multimodal dynamics, typically offering point forecasts rather than distributions, which carry richer information for decision-making. In contrast, diffusion forecasters like TimeGrad [4], can effectively guide multistage stochastic optimization by learning the full conditional distribution of future observations. Other diffusion-based models for time series, such as those designed for financial denoising [5] or integrated with foundation models [6], can similarly capture complex, temporal, multimodal dependencies. Diffusion-based methods have been applied to scenario generation in energy domains. For example, Li et al. [7] use a multi‐resolution diffusion model with time-series decomposition for renewable energy outputs, and Zhao et al. [8] propose a conditional diffusion model to jointly handle source-load uncertainties with spatio-temporal fusion, controllability and noise estimation. However, these works generate scenario sets rather than structured scenario trees, and do not explicitly integrate scenario trees into multistage decision optimization.
In this paper, we introduce the Diffusion Scenario Tree (DST) framework, a generic methodology for constructing scenario trees using a diffusion-based probabilistic forecaster. Each node represents a potential future state along with its associated realization probability, recursively generated from the learned distribution. This structure can be integrated naturally in optimization frameworks, providing the decision-maker with a rich representation of the forecast uncertainty.
We demonstrate the effectiveness of our approach by applying it to a multistage stochastic optimization problem for forecasting electricity prices in the New York power grid. We compare the performance of DST against scenario trees generated from VAR and LSTM models, our prior diffusion-based methods which showcased powerful capabilities, and Model-Free Reinforcement Learning (RL) baselines, showing superior behavior in optimizing decisions under uncertainty.
Our previous work [1] introduced Diffusion-Informed MPC, a framework that integrates diffusion forecasts into MPC algorithms. The present work advances our research by proposing a fundamentally new algorithm that generates diffusion-based scenario trees for multivariate prediction, tested on the same problem setup to enable direct comparison with standard MPC, and showcase the benefits of using DST.
2 Background
Decision-making in applications like energy markets, finance, and robotics requires precise modeling of time-evolving uncertainty to optimize actions. Scenario trees, a standard tool used in multistage stochastic optimization, represent uncertainty through nodes for possible future states and branches with associated probabilities [2, 3]. These trees respect non-anticipativity so that decisions at each stage depend only on observed information up to that point, and they discretize continuous distributions, enabling tractable optimization.
Early methods for scenario tree generation use sampling and scenario reduction [2, 9], while more recent approaches rely on Markov chain approximations or moment-matching to capture multivariate dynamics [3, 10]. Forecasting models are often used to generate scenarios across stages. Time series models like ARIMA or GARCH generate scenarios from historical data, followed by recursive reduction or bundling [11, 12]. Data-driven methods align scenarios with statistical properties such as mean and variance through optimization and clustering [13]. Non-parametric density estimation, like kernel methods, samples trajectories without fixed forms, capturing complex patterns [14], while parametric models such as SDEs for price dynamics assume structured distributions [15]. Beyond these, recent applications include learning-based adaptive scenario-tree MPC with Bayesian neural networks [16], scenario-tree MPC-based RL for autonomous vehicles [17], scenario-tree MPC for highway interactions [18], microgrid sizing with scenario trees [19].
However, these methods face challenges: statistical forecasters assume normal distributions for the errors, limiting their ability to model complex patterns; traditional forecasting models do not have the required inherent complexity to capture multimodal dynamics, and scenario reduction can compromise tree robustness by discarding useful scenarios.
In contrast, we propose to use diffusion-based forecasters like TimeGrad [4] to acquire accurate, probabilistic predictions. TimeGrad, is an autoregressive model that guides the diffusion processes of a Denoising Diffusion Probabilistic Model (DDPM) [20] using a Recurrent Neural Network to encode the time series history, learning the conditional distribution of future multivariate time series:
$p_{\theta}(\mathbf{x}_{k_{0}:N}|\mathbf{x}_{1:k_{0}-1})\approx q(\mathbf{x}_{k_{0}:N}|\mathbf{x}_{1:k_{0}-1}),\vskip-5.0pt$
where $\mathbf{x}_{k}\in\mathbb{R}^{D}$ is the multivariate time series vector at time $k$. TimeGrad has showcased impressive ability to capture long-range dependencies and multimodal patterns, making it ideal for sampling trajectories to construct scenario trees.
Our framework integrates TimeGrad’s probabilistic forecasts into scenario tree construction, using clustering to organize trajectories. This enables more efficient optimization, achieving greater accuracy and robustness compared to using VAR and LSTM, and over standard deterministic and stochastic MPC using the exact same TimeGrad model.
3 Methodology
3.1 Problem Definition
We present a method for a class of partially observable stochastic systems, where dynamics are separable into a deterministic component, influenced by actions, and a stochastic one, driven by external factors. This class of problems with separable state is common across many real-world applications such as energy market trading, finance, inventory management, and robotics. We model the system as a Partially Observable Markov Decision Process (POMDP), defined by $(S,A,T,R,\Omega,O,\gamma)$, where $S$ is the state space, $A$ the action space, $T$ the transition function, $R$ the reward function, $\Omega$ the observation space, $O$ the observation function, and $\gamma\in[0,1)$ the discount factor, consistent with our previous work [1].
At time step $k$, the environment is in a hidden state $s_{k}\in\mathcal{S}$. The agent selects an action $a_{k}\in\mathcal{A}$, causing the environment to transition to state $s_{k+1}$, yielding observation $o_{k+1}\sim O(o_{k+1}|s_{k+1},a_{k})$ and reward $r_{k}=R(o_{k},a_{k})$. The observation is $o_{k}=(o_{k}^{d},o_{k}^{s})\in\Omega$, where $o_{k}^{d}$ is the deterministic part (e.g., battery state-of-charge) evolving as $o_{k+1}^{d}=f(o_{k}^{d},a_{k})$, and $o_{k}^{s}\in\mathbb{R}^{D}$ the stochastic part (e.g., energy prices). In applications like energy bidding, true states are unmeasurable due to unpredictable factors (e.g., demand, weather, human factor), so the agent must rely on observations.
The agent’s objective at time $k_{0}$, given the observation history $o_{1},\dots,o_{k_{0}}$, is to find a policy $\pi$ maximizing the expected cumulative discounted reward over a finite horizon:
$\underset{\pi}{\text{maximize}}\quad\mathbb{E}\left[\sum_{k=k_{0}}^{k_{0}+N-1}\gamma^{k-k_{0}}R(o_{k}^{c},o_{k}^{u},a_{k})\right],\vskip-5.0pt$
which we propose to approximate with a finite-horizon multistage scenario tree-based MPC problem:
$\underset{\pi}{\text{maximize}}\quad\left[\sum_{k=k_{0}}^{k_{0}+N-1}\gamma^{k-k_{0}}R(o_{k}^{d},o_{k}^{s},a_{k})\right],\vskip-5.0pt$
where, the agent forecasts the stochastic part $\hat{o}_{k_{0}+1:k_{0}+N}^{s}$, organizing it into a scenario tree for tree-based optimization.
3.2 Diffusion-Based Forecasting
We use TimeGrad, a diffusion-based autoregressive model, to forecast the stochastic component’s conditional distribution $p_{\theta}(o_{k+1}^{s}|o_{1:k}^{s})$. At each decision epoch, TimeGrad generates $M$ sample forecasts, and organizes them in tree nodes.
3.3 Scenario Tree-Based Stochastic Optimization
Scenario Tree-based MPC optimizes decisions over a finite horizon using a scenario tree, with the root as the current observation $o_{k_{0}}$, nodes at each decision stage $t$ representing possible observation realizations, and branches are associated with probabilities of transitioning to each node [21]. It is important to note that the branching frequency may differ from the observation time step frequency (e.g., in the experiments discussed in Section 4 the branching step is daily while the observations are hourly). Therefore, a node in the scenario tree may represent a sequence of observations, all acquired at one decision stage. Non-anticipativity ensures decisions at stage $t$ depend solely on information available up to $t$, enforcing identical actions for scenarios sharing the same history. An example of such a scenario tree is shown in Figure 1.
$\mathbf{o}_{t}$$\mathbf{\hat{o}}_{t+1}^{(1)}$$\mathbf{\hat{o}}_{t+1}^{(2)}$$\mathbf{\hat{o}}_{t+2}^{(1)}$$\mathbf{\hat{o}}_{t+2}^{(2)}$$\mathbf{\hat{o}}_{t+2}^{(3)}$$\mathbf{\hat{o}}_{t+2}^{(4)}$$P(\mathbf{\hat{o}}_{t+1}^{(1)},w_{t+1}^{(1)})$ $a_{t}$ $P(\mathbf{\hat{o}}_{t+1}^{(2)},w_{t+1}^{(2)})$ $a_{t}$ $P(\mathbf{\hat{o}}_{t+2}^{(1)},w_{t+2}^{(1)})$ $a_{t+1}^{(1)}$ $P(\mathbf{\hat{o}}_{t+2}^{(2)},w_{t+2}^{(2)})$ $a_{t+1}^{(1)}$ $P(\mathbf{\hat{o}}_{t+2}^{(3)},w_{t+2}^{(3)})$ $a_{t+1}^{(2)}$ $P(\mathbf{\hat{o}}_{t+2}^{(4)},w_{t+2}^{(4)})$ $a_{t+1}^{(2)}$ Realized observation at stage $t$ Realized observation at stage $t+1$ Realized observation at stage $t+2$
Fig. 1: Visualization of a Scenario Tree that discretizes the probability distribution of the future steps.
3.3.1 Diffusion Scenario Tree
The Diffusion Scenario Tree method, described in Algorithm 1, constructs the tree, stage-by-stage, from the root using TimeGrad. For stage $t$, at each node, TimeGrad samples $M$ forecasts $\{\hat{o}^{s,(i)}_{t}\}_{i=1}^{M}\subseteq\mathbb{R}^{H\times D}$ for the stochastic component, from the conditional distribution of future timesteps, conditioned on past observations from the node’s predecessors, over the stage horizon $H$. These forecasts are clustered using K-means into $K$ groups, with centroids serving as representative predictions for the nodes. Branch probabilities are computed as $p_{k}=\frac{1}{M}\sum_{i=1}^{M}\mathds{1}(\ell_{i}=k)$, and the top-$L$ child nodes are retained based on the highest cumulative probabilities of the paths from the root to these nodes at stage $t$. The retained nodes’ probabilities are normalized to sum to 1. Each cluster centroid is used as the input for forecasting at the next stage. Therefore DST recursively samples future trajectories and generates nodes based only on tree paths, ensuring the non-anticipativity constraint (decisions depend only on observed history) by construction.
3.3.2 Solving the Multistage Stochastic MPC
Given a scenario tree $\mathcal{T}$, we solve the multistage stochastic MPC to optimize decisions across $N$ stages:
$\underset{\{a^{i}_{t}\}}{\text{maximize}}$ $\quad\sum_{t=k_{0}}^{k_{0}+N-1}\sum_{i\in\mathcal{T}_{t}}\pi_{i}R(o^{d,i}_{t},o^{s,i}_{t},a^{i}_{t})$
subj. to $o^{d,1}_{k_{0}}=o^{d}_{k_{0}},\quad o^{s,1}_{k_{0}}=o^{s}_{k_{0}},$
[…]
$a^{i}_{t}=a^{j}_{t}\;\forall i,j\in\mathcal{T}_{t}\text{ for same history},a^{i}_{t}\in\mathcal{A}\;\forall i,t.$
Here, $\mathcal{T}_{t}$ denotes nodes at stage $t$, with probability $\pi_{i}$, observation $(o^{d,i}_{t},o^{s,i}_{t})$, and decision $a^{i}_{t}$. The deterministic component evolves via $f$, while the stochastic part is forecasted.
4 Results
We evaluate the Diffusion Scenario Tree (DST) framework on the energy arbitrage task in New York’s day-ahead electricity market, where a Battery Energy Storage System (BESS) optimizes energy transactions based on forecasted prices. The environment is modeled as a POMDP, with observations consisting of battery state-of-charge (SoC) and electricity prices, actions control battery charge, and rewards defined by the trading revenue minus degradation costs, defined thoroughly in [1]. In this problem, the deterministic component, battery State of Charge (SoC), evolves via known dynamics $f$, and the stochastic component’s (energy prices) dynamics are estimated probabilistically by TimeGrad. Experiments use 10 months of hourly electricity price data, covering multiple seasons, to test on diverse market conditions [^1] Information about the implementation of the models: https://sites.google.com/view/stelios-zarifis-thesis-ntua/, in Section ”Experimental Results”..
Figure 2 shows the optimized actions by Stochastic MPC based on our DST over a 3-day horizon. Dotted lines show forecasted prices in the scenario tree (branching daily), the gray line actual prices published after optimization, the blue line shows the applied optimal actions for the first stage, and dashed lines show how SoC evolves when applying the branch actions. DST SMPC applies the actions for the first stage and then re-plans, as new observations are available. Visually, we can see that the agent buys and sells energy at near-optimal points, generating profit from price differences across the horizon.
Fig. 2: Strategy planned by Scenario tree-based MPC. Dotted lines depict the forecasted stochastic component, organized in a tree. Gray line shows the actual stochastic part evolution, blue line represents the optimized actions for the first stage, and dashed lines the optimized actions across branches.
4.1 DST-based MPC vs. Standard MPC
Table 1 compares DST-guided SMPC with benchmarks: Perfect MPC (full knowledge of the realized trajectories), Oracle MPC (perfect forecasts over the same finite horizon), and Monte Carlo SMPC (sampling multiple trajectories from the model’s distribution to estimate the expectation of the objective via Monte Carlo averaging, with its number of realized trajectories equal to DST’s scenarios), the best-performing MPC using the diffusion-based forecaster from our previous work [1], which also outperformed all implementations relying on conventional forecasters. DST SMPC achieves an average reward of $85.81$, outperforming MC SMPC. This slight improvement is significant for several reasons. First, Monte Carlo SMPC is already a powerful model, achieving near-optimal performance. Second, the prediction horizon in our setup is relatively long, meaning that the influence of early-stage decisions diminishes over time. While later-stage decisions still impact the overall performance, their effect is less pronounced. Third, DST SMPC adheres to non-anticipativity, ensuring that decisions at each stage depend solely on information available up to that point. This aligns with real-world decision-making scenarios, where future information is not accessible. In contrast, standard MPC approaches may lead to decisions that are not as realistic in practice.
4.2 DST-based MPC vs. LSTM- and VAR-based MPC
Table 1 also shows DST SMPC versus LSTM- and VAR-based scenario tree SMPC. Our approach, surpasses both LSTM-based and VAR-based implementations by 18.0% and 25.4%, respectively. Consistently superior across all tests, DST’s ability to discretize multimodal distributions provides a more accurate model to guide the MPC. Another point highlighting the benefit of using scenario tree-based MPC over standard MPC is that even the classical forecasters (LSTM, VAR), when used recursively in a scenario tree manner, provide more accurate actions, as multiple scenarios are considered instead of a single one in standard MPC. This leads to an improvement of approximately 91.2% for the LSTM-based MPC and 10.7% for the VAR-based MPC over the experiments of our previous work [1].
4.3 DST-based MPC vs. Model-Free RL
Table 2 compares DST SMPC with a MFRL Deep Q Network (DQN) (state consisting of 12 future prices). Our approach exceeds the best DQN implementation by 218.8%, highlighting the advantage of using a dynamics model, when it exists.
Input: $\mathcal{T}$: empty tree; $\mathcal{O}$: observation history; $D$: tree depth;
predictor: diffusion-based forecaster; $K$: clusters per stage; $M$: forecast samples; $H$: stage horizon; $L$: top-$L$ children.
Output: Scenario tree with forecasts & probabilities
Init. queue $Q\leftarrow\{(\text{root},\mathcal{O},0)\}$, node_counter = 0;
while $Q\neq\emptyset$ do
 // node, obs history, stage
 $(node,\mathcal{O},\ell)\leftarrow\text{Dequeue}(Q)$
 $node.\text{id}\!\leftarrow\!\text{node\_counter}$++ // Assign unique ID
 if $\ell\geq D$ then
    continue // maximum depth reached
 // Sample conditioned to history
 $\{\mathcal{F}^{(i)}\}_{i=1}^{M}=\texttt{predictor.sample}(\mathcal{O}^{s},M)$
 // Cluster in groups
   Reshape forecasts into $X\in\mathbb{R}^{M\times(H\cdot D)}$
  Cluster $X$ with K-means $\to$ labels $\{\ell_{i}\}$
 for $k=1,\dots,K$ do
    // Representative forecast, branch & cumulative prob
    $\mu_{k}\leftarrow\text{centroid of cluster }k$
    $p_{k}\leftarrow\frac{1}{M}\sum_{i=1}^{M}\mathds{1}(\ell_{i}=k)$
    $P_{k}\leftarrow node.P\cdot p_{k}$
     Create child node $v_{k}$ with $(\mu_{k},P_{k})$;
  Sort $\{v_{k}\}$ by $P_{k}$, keep top $L$ ;
 // Prune tree
  Normalize probabilities: $P_{k}\leftarrow P_{k}/\sum_{j=1}^{L}P_{j}$
 for each kept $v_{k}$ do
    $\mathcal{O}^{\prime}\leftarrow(\mathcal{O}\,||\,\mu_{k})$ // Update obs history
     Attach child $v_{k}$ to nodenode
[…]
return $\mathcal{T}$ // Final scenario tree
Algorithm 1 Diffusion Scenario Tree
5 Conclusion
We proposed the Diffusion Scenario Tree (DST) framework, which constructs scenario trees using diffusion-based probabilistic forecasters like TimeGrad to improve multistage stochastic optimization. DST recursively samples trajectories, to expand the scenario tree, and organizes them via clustering, ensuring non-anticipativity. The resulting structure not only enables better decision-making but also provides interpretability, as each branch clearly represents a possible future scenario and the corresponding optimal actions. Evaluated on a real-world decision problem, our method outperforms the same scenario tree algorithms guided by more conventional predictive models, the standard MPC implementations guided by diffusion-based forecasters, and RL baselines.
Future work includes problem-driven scenario tree generation, so that the generated scenarios align with the problem’s structural characteristics and constraints [22, 23].
Table 1: Comparison: DST-based vs Standard, LSTM-, and VAR-based MPCs (higher is better).
Month Perfect Oracle MC DST LSTM ST VAR ST
Month MPC MPC SMPC SMPC SMPC SMPC
2018-06 $99.98$ $97.13$ $\underline{90.05}$ $\mathbf{94.47}$ $87.94$ $60.51$
2018-07 $147.46$ $141.56$ $\underline{134.87}$ $\mathbf{140.00}$ $116.23$ $107.92$
2018-08 $123.11$ $115.31$ $\underline{111.28}$ $\mathbf{112.63}$ $109.21$ $97.11$
2018-09 $108.44$ $102.32$ $\underline{99.67}$ $\mathbf{100.20}$ $96.09$ $79.63$
2018-10 $104.05$ $95.36$ $\mathbf{102.13}$ $\underline{101.17}$ $81.43$ $70.20$
2019-04 $60.35$ $54.62$ $\mathbf{52.19}$ $\underline{45.00}$ $36.14$ $34.25$
2020-12 $68.85$ $59.31$ $\underline{57.41}$ $\mathbf{59.92}$ $54.37$ $55.97$
2021-01 $64.03$ $59.00$ $\underline{55.87}$ $\mathbf{58.27}$ $52.56$ $51.84$
2021-02 $109.34$ $108.33$ $\mathbf{98.20}$ $\underline{95.71}$ $50.99$ $78.03$
2021-03 $56.94$ $56.35$ $\mathbf{53.95}$ $\underline{50.72}$ $42.37$ $49.12$
Sum $942.57$ $889.30$ $\underline{855.63}$ $\mathbf{858.09}$ $727.32$ $684.58$
Average $94.26$ $88.93$ $\underline{85.56}$ $\mathbf{85.81}$ $72.73$ $68.46$
Table 2: DST SMPC vs MFRL in 10 months (higher is better).
Model Sum Average
DQN ($32\times 64\times 32$) $55.76$ $5.58$
DQN ($64\times 128\times 64$) $\underline{269.15}$ $\underline{26.92}$
DQN ($128\times 256\times 128$) $151.80$ $15.18$
DQN ($64\times 128\times 128\times 64$) $51.02$ $5.10$
DQN ($64\times 128\times 256\times 128\times 64$) $196.27$ $19.63$
DST SMPC $\mathbf{858.09}$ $\mathbf{85.81}$
6 Compliance with Ethical Standards
This is a numerical simulation study for which no ethical approval was required.